LBC/LUCI Technical Observing, prep for Partner Science:
UT 2011 Jan 29

Observer: OKuhn (Tucson), DThompson (LBT)
Telescope Operator: DHuerta
Instrument Support: JMorris
Telescope Support:
SW support:

Summary

00:06 Toutside 3.3C, Tmirror: SX=-0.7 DX=-0.7. RH=4.0%. Clouds low on S horizon.

  • B/R twilight flats at 0, 90, 180 degrees PA
  • Checked co-pointed range-balanced collimation models...it works well!
  • Lost 1 hr to LBCB restarts (bad temperature readings)
  • Checked filter background count rates and filter positioning
  • Guiding and closed-loop focus tests on M82
  • Rotator center updated on both cameras
  • Pointing observations: 32 points
  • Filter focus offsets remeasured (took 2 full passes of data)
  • Additional guide tests, NGC5713 field
  • 900s darks in both cameras

Almanac:
  • sunset at 17:58 ---- sunrise at 7:07 MST
  • Naut. (12d) twilight: 18:44 --> 6:21 MST
  • Astron. (18d) twilight: 19:13 --> 5:52 MST
  • Time between astron. twilights: 10.6 hours (13.2h between sunsets)
  • LST @ midnight = 08:13:10.7

Details

00:32 BIASDARK/25Bias_Bino.ob
00:48 BIASDARK/2Darks_Bino.ob skipped, not enough time (done at end of night).

01:10 Taking twilight flats...
  • 01:18 SkyFlat_BR_pa0_5.ob scaled by 1.5 (12k, 14k at end)
  • 01:22 SkyFlat_BR_pa180_5.ob scaled by 3.0 (9k, 10k at end)
  • 01:27 SkyFlat_BR_pa90_5.ob scaled by 12.0 (11k and 12k at end)

Image timestamps look OK.

01:40 Changed to new range-balanced co-pointed collimation models

01:42 ACT0295 DOFPIA...no good pupils!
01:50 ACT0280 DOFPIA...pupils too large. SX +16000 Z4, DX +20000 Z4
01:59 ACT0280 DOFPIA...exit after iteration 16. Things are close, but affected by warm mirrors (+3, +4 deg rel).

02:26 RB-fastextra4.ob takes all 4 chips w/additional DX Z4 of -20000nm to make the pupils larger. The dofpia pupils were looking a little odd.

02:37 ACT0280 do lbcrangebal here, EL~75 B23750 R23743, Confirm: B24047 R24047 B@1040, 2920 R@1080,2918

02:41 ACT0210 Check co-pointing EL~57 B24359 R24353, lbcrangebal, Confirm: Bxxxxx R24723 (CCD Controller error!)
2109   2011/01/29   02:47:36.383379         camera system encountered a problem
2108   2011/01/29   02:47:36.383288   CAMERA       CCD Controller error errcode:9 [src/camera/camera.c:1099]
2107   2011/01/29   02:47:36.383225   CAMERA       reinitializing CCD Controller failed 
Turn off and then turn on again seems to have recovered.

02:52 Confirm again at EL~57... B25254 R25250

02:54 ACT0123 Check co-pointing at EL~32 B25523 R25517, lbcrangebal, confirm: B25803 R25759. Co-pointing corrections were small (~few arcsec) at each EL once corrected initially.

03:00 FILTERtesting/testfiltersBR_onsky.ob 5h+40d, EL~79

03:20 Run dofpia here, EL~80. Blue CCD Controller error again, at iteration 7. Turn off and then turn on...03:36

3577   2011/01/29   03:34:00.622000   CAMERA    CCDCTRL    temperature safety limit exceded (-47.6Celsius instead of -50.0Celsius)
3576   2011/01/29   03:34:00.622000   CAMERA    CCDCTRL    detectors temperature: 225.6 [Kelvin] 

...but BLUE encountered the same temperature problem...
77   2011/01/29   03:39:49.415850         camera system encountered a problem
76   2011/01/29   03:39:42.142340   CAMERA       CCD Controller error, errcode:9 [src/camera/camera.c:556]
75   2011/01/29   03:39:42.055000   CAMERA    CCDCTRL    safetyoff completed
74   2011/01/29   03:39:42.040000   CAMERA    CCDCTRL    safetyoff ...
73   2011/01/29   03:39:42.024000   CAMERA    CCDCTRL    temperature safety limit exceded (-34.7Celsius instead of -50.0Celsius)
72   2011/01/29   03:39:42.008000   CAMERA    CCDCTRL    detectors temperature: 238.4 [Kelvin] 

...turn off and on again...

54   2011/01/29   03:42:48.091477         camera system encountered a problem
53   2011/01/29   03:42:40.797371   CAMERA       CCD Controller error, errcode:9 [src/camera/camera.c:556]
52   2011/01/29   03:42:40.751000   CAMERA    CCDCTRL    safetyoff completed
51   2011/01/29   03:42:40.735000   CAMERA    CCDCTRL    safetyoff ...
50   2011/01/29   03:42:40.719000   CAMERA    CCDCTRL    temperature safety limit exceded (-32.2Celsius instead of -50.0Celsius)
49   2011/01/29   03:42:40.704000   CAMERA    CCDCTRL    detectors temperature: 241.0 [Kelvin]
48   2011/01/29   03:42:38.488000   TRACKERS   CCDCTRL    load FPGA on SPC-A200 failed
47   2011/01/29   03:42:14.740081   ROTATOR       GOYA no prompt return [src/rotator/rotator.c:2078]
46   2011/01/29   03:42:13.420064   ROTATOR       GOYA no prompt return [src/rotator/rotator.c:2078]
45   2011/01/29   03:39:49.415850         camera system encountered a problem
44   2011/01/29   03:39:42.142340   CAMERA       CCD Controller error, errcode:9 [src/camera/camera.c:556] 

03:45 Off and on again...and again the same error:

20   2011/01/29   03:46:29.575000   CAMERA    CCDCTRL    temperature safety limit exceded (-36.0Celsius instead of -50.0Celsius)
19   2011/01/29   03:46:29.559000   CAMERA    CCDCTRL    detectors temperature: 237.2 [Kelvin] 

03:48 Turn off (also HK) and then on again...worked this time!

03:55 DOFPIA here, EL~82

51   2011/01/29   03:58:35.731241         OB execution error: none [src/lbc/actions.c:1639]
50   2011/01/29   03:58:35.611533         >>> preset/expose/save interrupted at expose time
49   2011/01/29   03:58:34.995000   TRACKERS   CCDCTRL    safetyoff completed
48   2011/01/29   03:58:34.979000   TRACKERS   CCDCTRL    safetyoff ...
47   2011/01/29   03:58:34.948000   TRACKERS   CCDCTRL    safetyoff completed
46   2011/01/29   03:58:34.948000   TRACKERS   CCDCTRL    safetyoff ...
45   2011/01/29   03:58:34.776968         camera system encountered a problem
44   2011/01/29   03:58:34.776888   CAMERA       CCD Controller error errcode:9 [src/camera/camera.c:1185]
43   2011/01/29   03:58:34.776852   CAMERA       reinitializing CCD Controller failed
42   2011/01/29   03:58:34.719000   CAMERA    CCDCTRL    safetyoff completed
41   2011/01/29   03:58:34.704000   CAMERA    CCDCTRL    safetyoff ...
40   2011/01/29   03:58:34.688000   CAMERA    CCDCTRL    temperature safety limit exceded (-35.7Celsius instead of -50.0Celsius)
39   2011/01/29   03:58:34.672000   CAMERA    CCDCTRL    detectors temperature: 237.4 [Kelvin] 

04:01 Off and on again...and failed again...
45   2011/01/29   04:02:13.572000   CAMERA    CCDCTRL    temperature safety limit exceded (-36.7Celsius instead of -50.0Celsius)
44   2011/01/29   04:02:13.556000   CAMERA    CCDCTRL    detectors temperature: 236.5 [Kelvin] 

04:05 Off and on again...and failed again...
43   2011/01/29   04:05:57.962000   CAMERA    CCDCTRL    temperature safety limit exceded (-38.8Celsius instead of -50.0Celsius)
42   2011/01/29   04:05:57.946000   CAMERA    CCDCTRL    detectors temperature: 234.4 [Kelvin] 

04:07 Moving to horizon, JMorris to check controllers, fans. He reports all fans on and all boxes feel cold to the touch. Cycling power again, including HK. Now blue trackers taking longer than usual to come up. There is a warning message in the log:
3   2011/01/29   04:26:28.749000   TRACKERS   CCDCTRL    load FPGA on SPC-A200 failed
2   2011/01/29   04:25:55.588000   TRACKERS   CCDCTRL    load FPGA on SPC-A200 failed 

but the system seems to be back up at 04:31. 1 hour lost to LBC Blue

04:31 Preset back to testfiltersBR_onsky, abort and the DOFPIA here to collimate

SX 0.000 0.000 0.283 -22.632 3.055 0.000 DX 0.000 0.000 0.047 -4.099 6.043 0.000

04:46 ROTCEN check B44659 R44655 at PA=0, B44853 R44850 at PA=180, B45103 R45101 at PA=0 at EL~75
RED: Old rotator center: 1080. 2912.
New rotator center:
Xc=1078.85 Yc=2912.91 from geomap.
Xc=1078.62 Yc=2913.02 from averages.

BLUE Old rotator center: 1040. 2921.
New rotator center:
Xc=1035.63 Yc=2923.34 from geomap.
Xc=1034.98 Yc=2923.47 from averages.

04:53 ACT0331 to collimate, then rangebal, EL=84.6. Red pupuils have been looking oddly not-round tonight. Converged, but it took 17 iterations! Olga is updating the LBC headers with the new rotator centers although they are not far off. Running DOFPIA again just to stay collimated. And now even though BLUE is collimated, there is a noticeable bright ring around the outer edge of the pupils.

05:30 Last blue collimation sent -1000 Z11 and -100 Z22. Clearing all on SX back to the collimation model and re-run DOFPIA. Blue failed because of coma, added +2000nm Z7 and rerun dofpia. 05:41 converged, and Olga is ready so I am stopping the DOFPIA.

05:42 ACT0331 (same field) for lbcrangebal at EL~79. Confirmed that the headers have been changed to the new rotcen values.
lbcrangebal sent this for the instrument offsets (confirmed on PSFGUI)
SX: dX = 0.145 dY = 0.049 dRX = 1.049 dRY = -3.118
DX: dX = -0.072 dY = 0.143 dRX = 3.074 dRY = 1.539

05:51 ACT0161 at EL~45. Preset timed out because of slow AZ rotation. Resent at 05:58 B55816 R55811
SX: -0.001 0.101 0.070 2.163 0.025 0.000 (X Y Z RX RY RZ)
DX: 0.095 0.117 0.079 2.511 -2.030 0.000

06:01 ACT0095 at EL~24. B60322 R
1574   2011/01/29   06:04:55.603000   CAMERA    CCDCTRL    hardware problems [ccdctrl.c:1373]
1573   2011/01/29   06:04:55.603000   CAMERA    CCDCTRL    "CCD Controller fiber link error or incomplete data" (code:22) controller error [ccdctrl.c:2378] 
But LBC seems to be trying to restart the red camera automatically. Nope...CCD Controller Error. Turn off and then on again. Nope:
27   2011/01/29   06:10:11.150540         camera system encountered a problem
26   2011/01/29   06:10:02.219936   CAMERA       RPC Client/Host communication failed retval:5 [src/camera/camera.c:530] 

Turn off and then on again...OK this time.

06:15 ACT0095 at EL~23 B61614 R61610
Blue is a bit out of focus, but no strong coma. SX: 0.071 0.086 0.070 1.841 -1.524 0.000
DX: 0.058 0.150 0.079 3.221 -1.243 0.000

06:25 ACT0404 (near M82) for DOFPIA and rangebal. B63954 R63949
SX: 0.113 0.023 0.070 0.504 -2.425 0.000
DX: -0.086 0.135 0.079 2.886 1.850 0.000

06:43 M82_BRSY_focus.ob (~10' north of M82 proper to avoid extended source). This is to test guiding, focus control, and dithering. DOFPIA stopped after 7 iterations. It was close, but not quite converged. Close enough for these tests.

06:55 M82_BRSY_Rmaster.ob (2.9...5 pixels FWHM across all bands. Guiding seems to have worked well. Point sources stayed compact and round throughout the sequence.
07:23 M82_BRSY_focus.ob to re-run dofpia.
07:33 M82_BRSY_Bmaster.ob set up with shorter exposures. R still saturating on a few of the brightest sources in the core, others are all OK.

08:24 ACT0392 in the NE (AZ > 360), EL~54. DOFPIA, lbcrangebal, then start a pointing run. Close, but only 2 pupils in blue, 3 in red. Moving to WT10_350 (first pointing star).

08:37 WT10_350 EL~73 (AZ 393). Blue seems to be oscillating in Z5/Z6 just above the threshold. Converged...lbcrangebal B84759 R84754. SX: -0.075 0.009 0.070 0.197 1.607
DX: 0.080 0.153 0.079 3.273 -1.713

Pointing Model Observations (See ~telescope/PointingLogs/20110129.log)
N SourceID B R
01 WT10_350 85116 85112
02 ACT0403 85417 85412
03 ACT0402 85644 85639
04 WT10_348 85931 85926
05 WT10_299 90240 90234
06 WT10_199 90548 90543
07 ACT0237 90843 90839
08 ACT0084 91146 91141
09 WT10_152 91430 91426
10 ACT0275 91714 91709
11 ACT0209 92002 91956
12 WT10_341 92258 92253
13 WT10_326 92501 92458
14 ACT0296 92712 92707
15 ACT0262 92921 92916
16 ACT0307 93217 93212
17 ACT0169 93531 93526
18 ACT0086 93823 93815
19 ACT0116 94027 94025
20 WT10_290 94322 94317
21 WT10_254 94601 94556
22 WT10_233 94849 94844
23 ACT0286 95106 95101
24 WT10_322 95408 95403
25 ACT0404 95723 95719
26 WT10_400 95951 95946
27 WT10_406 100208 100202
28 WT10_337 100523 100523
29 WT10_376 100814 100809
30 ACT0380 101117 101112
31 WT10_401 101345 101340
32 WT10_407 101622 101618

10:18 ACT0241 for Filter focus offsets, EL~67. Start times for each OB are listed below. Dofpia was run before each block of three.
  • 10:38, 11:39 RB_zB_superfoc.ob
  • 10:43, 11:43 RB_rV_superfoc.ob
  • 10:47, 11:47 RB_YU_superfoc.ob

  • 10:58, 11:58 RB_Ir_superfoc.ob
  • 11:03, 12:02 RB_rV_superfoc.ob
  • 11:07, 12:06 RB_RUs_superfoc.ob

  • 11:16, 12:15 RB_ig_superfoc.ob
  • 11:20, 12:18 RB_rV_superfoc.ob
  • 11:24, 12:22 RB_zB_superfoc.ob

11:32 Lost tracking, not obvious why. Resend preset to ACT0241, dofpia, then loop back up above to record times of second set of focus data.

12:26 Check exposure scaling on NGC 5713. Took some dithered BR data here, all 40s B/R

13:20 Close dome. When parked, run 2Darks_Bino.ob





LBC Observing Plan

OBs in ~LBTO/Calib_OBs/...
Warmup for 15min before turn on
Health and stability checks:
  • BIASDARK/25Bias_Bino.ob (dome lights off!)
  • BIASDARK/2Darks_Bino.ob.
Check image timestamps
Twilight flats (Mkskyflat/)
  • PA = 0,180 to check for differences
  • We have BR + SY, do V+I for M82?
  • Do filters to be used in PR/sci test OBs: Us,B,g,r Y, I, i combinations:
    • UsY (1:0.3 exptime ratio) look in /home/LBTO/Calib_OBs/SKYFLAT_TEST_SCALING/
    • BI (1:0.3 exptime ratio) in /home/LBTO/Calib_OBs/SKYFLAT/
    • gi (1:1)
    • rr (1:1)
  • note scalings for adjustment of OB exposure times.
Install new co-pointed collimation models from Andrew
Check the filters vs. sky counts near zenith
  • Also check filter alignment/vignetting
  • OB: FILTERtesting/testfiltersBR_onsky.ob
  • Edit TEC_RA and TEC_DEC to appropriate values. Now TEC_RA=75 and TEC_DEC=70).
Astrometry - obtain one astrometric field.
  • OBs: /home/lbcobs/REDCOMMISSIONING/ASTROMETRIC_STD/
  • Copy these OBs to ~LBTO/Calib_OBs
Throughput - obtain one standard star field if photometric.
  • OBs: STANDARDS/
Astronomical test on M82?
  • Extended source - where to do DOFPIA?
  • Check guiding, closed-loop focus control, dithering
  • Do V+I to go with the twilight flats

Filter focus offsets
  • OBs: SUPERFOC/...
  • RB_zB_superfoc.ob
  • RB_rV_superfoc.ob
  • RB_YU_superfoc.ob

  • RB_Ir_superfoc.ob
  • RB_rV_superfoc.ob
  • RB_RUs_superfoc.ob

  • RB_ig_superfoc.ob
  • RB_rV_superfoc.ob
  • RB_zB_superfoc.ob

Work on Blue convergence
  • Need blue matrix separated from red
Pointing data with co-pointed collimation model

LUCI Observing Plan

Restart:
  • TelescopeService (build switch to BP7)
  • InstrumentManager (to correct lamp indicator problem)

LUCI1 functional checkout:
  • Check system time is same as our GPS ntp server
  • Run /home/engineer/bin/showSupervisorStatus.sh
  • Check LUCI1 Web-IO Control, verify that ROE and MCE are powered
  • Start LUCIFER Control SW and check that all is running as needed
  • Open UI and check that telescope data are being updated
  • Check darks (see: LUCIFER$/readnoise script)
    • O2DCR 2.0s (usually around 11.57 e-)
    • MER10 10.0s
  • Take sieve mask images in N3.75 and N1.8 to check focus gradients
  • Move calibration unit in (restart InstrumentManager first!)
    • Check that lamps work
    • Check gratings/tilts
    • Take an imaging flat and check levels
    • Move calibration unit out
  • Check mask tables after exchange
  • LUCI1 flexure tests (Nancy's script)

Correct initial collimation/pointing

Open dome SW tests, if requested
  • Optical pointing/co-pointing/range balancing algorithm, also using IRTC

Run science-type dice5 pattern to check that things still work (use M82?)
  • Preset to available BS91XX star from UI/Catalog
  • Execute script: ~/LBTO/SciencePrep/dice5_10x (~1 hr)
  • Calculate rms/position, scale & rotation of pattern

Repeat check of dithering along slit (20x 2' dither)
  • Preset to available BS91XX star from UI/Catalog
  • Execute script: ~/LBTO/SciencePrep/slitdither.acq
  • Align BS source to center of 2" longslit
  • Execute script: ~/LBTO/SciencePrep/slitdither.sci (~30 min)
  • Calculate AGw alignment to slit

Test focus task and verify it does what is needed on LUCI.

Repeat measurement of focal plane scale, orientation

Additional work on Andrew's M1/M2 collimation?

Take a set of (monocular) range-balanced collimation data
  • Stable conditions!
  • implement this collimation model

Collect field aberration data

Photometric zero point and extinction monitoring

Collect a "movie" of 2s LUCIFER images while an in-field offset is taking place.
  • Check to make sure offsets still happen in a reasonable manner

Collect through-focus sequence for Andrew - reasonably dense star field.
  • Try Z4 of -1000, -500. 0, 500, 1000 nm
  • Also check the AGw-LUCIFER focus in cold weather for temp term?

Collect astrometric field data to check AGw calibration for monitoring

Improve map of AGw patrol field limits at high field angle
  • Requires set up work on a stone field

Gather NB illumination correction data (BrG done on 20100606)

Re-test updatePointingReference functionality.

Continue gaining experience with LUCIFER on-sky (e.g. spec acq.)

Test dithering between extreme slits on MOS mask + telluric
  • Manual needs 0.59836mm/" scale for the dXdY calculation
  • The how-to set up telluric observations needs updating for clarity.

Investigate variable imaging FF gradients by taking data while the telescope is offset with respect to the enclosure (will need manual overriding of normal operations) to see if the observed gradients could be caused by something 'external to the instrument'.

-- OlgaKuhn, DavidThompson - 28 Jan 2011
Topic revision: r7 - 31 Jan 2011, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback